# Saber: An Efficient Sampling with Adaptive Acceleration and Backtracking Enhanced Remasking for Diffusion Language Model

### v1

@@ 1 Introduction
Diffusion language models (DLMs) have emerged as a promising non-autoregressive alternative in natural language processing (NLP) fields, with inherent advantages in parallel decoding and bidirectional context modeling through iterative denoising processes (Austin et al., 2021a, Ou et al., 2025, Nie et al., 2025, Ye et al., 2025b). Unlike existing autoregressive models (ARMs) that generate text left-to-right (Radford & Narasimhan, 2018, Radford et al., 2019, Brown et al., 2020, Touvron et al., 2023), DLMs can simultaneously refine multiple token positions by progressively unmasking the generation sequence, enabling global planning and iterative refinement (Ye et al., 2025a, Gong et al., 2025). This paradigm is especially compelling for the structured generation tasks like code generation.
@@ 3.1 Diffusion Language Models for Code
The current landscape of language models is dominated by the autoregressive paradigm (Radford & Narasimhan, 2018, Brown et al., 2020, Touvron et al., 2023, Dubey et al., 2024, Guo et al., 2025). However, their strict left-to-right and token-by-token generation process creates a major bottleneck for inference efficiency and inherently limits parallelism (Li et al., 2025). Therefore, a growing body of research for DLMs has emerged (Li et al., 2022a, Austin et al., 2021a, He et al., 2022), which operate through parallel generation and bidirectional context modeling to address the aforementioned constraints. Recently, large-scale DLMs, such as Dream (Ye et al., 2025b), DiffuLLaMA (Gong et al., 2024), and LLaDA (Nie et al., 2025) have demonstrated performance comparable to similar-scale ARMs, making them a highly promising alternative.
@@ D.2 Promising Architecture for Language Modeling
While the Transformer has been the foundational architecture for modern language models (Vaswani et al., 2017), the field is experiencing a significant shift with the rise of new paradigms (Dong et al., 2024c; 2025b). Mamba (Gu & Dao, 2023), leveraging a selective State Space Model, presents a compelling alternative that scales linearly with sequence length, effectively overcoming the quadratic complexity bottleneck of Transformers in long-context scenarios (Gu et al., 2021). Simultaneously, a fundamentally different approach is being explored with Diffusion models, which move away from traditional autoregressive generation (Li et al., 2022a). By learning to denoise a sequence from a random state, these models offer a unique framework for highly controllable and iterative text synthesis, signaling a potential new direction for generative AI.

### v2

@@ 1 Introduction
Diffusion language models (DLMs) have emerged as a promising non-autoregressive alternative in natural language processing (NLP) fields, with inherent advantages in parallel decoding and bidirectional context modeling through iterative denoising processes (Austin et al., 2021a, Ou et al., 2025, Nie et al., 2025, Ye et al., 2025b). Unlike existing autoregressive models (ARMs) that generate text left-to-right (Radford and Narasimhan, 2018, Radford et al., 2019, Brown et al., 2020, Touvron et al., 2023), DLMs can simultaneously refine multiple token positions by progressively unmasking the generation sequence, enabling global planning and iterative refinement (Ye et al., 2025a, Gong et al., 2025). This paradigm is especially compelling for the structured generation tasks like code generation.
@@ 3.1 Diffusion Language Models for Code
The current landscape of language models is dominated by the autoregressive paradigm (Radford and Narasimhan, 2018, Brown et al., 2020, Touvron et al., 2023, Dubey et al., 2024, Guo et al., 2025). However, their strict left-to-right and token-by-token generation process creates a major bottleneck for inference efficiency and inherently limits parallelism (Li et al., 2025). Therefore, a growing body of research for DLMs has emerged (Li et al., 2022a, Austin et al., 2021a, He et al., 2022), which operate through parallel generation and bidirectional context modeling to address the aforementioned constraints. Recently, large-scale DLMs, such as Dream (Ye et al., 2025b), DiffuLLaMA (Gong et al., 2024), and LLaDA (Nie et al., 2025) have demonstrated performance comparable to similar-scale ARMs, making them a highly promising alternative.
@@ D.2 Promising Architecture for Language Modeling
While the Transformer has been the foundational architecture for modern language models (Vaswani et al., 2017), the field is experiencing a significant shift with the rise of new paradigms (Dong et al., 2024c; 2025b). Mamba (Gu and Dao, 2023), leveraging a selective State Space Model, presents a compelling alternative that scales linearly with sequence length, effectively overcoming the quadratic complexity bottleneck of Transformers in long-context scenarios (Gu et al., 2021). Simultaneously, a fundamentally different approach is being explored with Diffusion models, which move away from traditional autoregressive generation (Li et al., 2022a). By learning to denoise a sequence from a random state, these models offer a unique framework for highly controllable and iterative text synthesis, signaling a potential new direction for generative AI.